A student has a weight of 655 N. While riding a roller coaster they seem
to weigh 1.96x103 N at the bottom of a dip that has a radius of 18.0 m. What
is the speed of the roller coaster at this point?
1
Expert's answer
2020-04-16T09:07:01-0400

Explanation

  • For the question to be meaningful, the value mentioned ("\color{red}\footnotesize\bold{1.96\times103\,N}" ) should be corrected to "\color{red}\footnotesize\bold{1.96\times10^3\,N}".
  • 655N weight happens to be when he is at rest.
  • The phenomenon of apparent weight occurs when there is a net force acting on the body (any acceleration, deceleration, upthrust etc).
  • What happens here is, when he take the path at the bottom of the dip he enters into a rotational motion in which a generated centrifugal force pushes him down towards the bottom with an intention to getting supplied the needed centripetal force to complete the motion.
  • There he feels as if his weight is increased.
  • We calculate the weight of an object by measuring the resultant force( as in the picture 1)
  • So you can see that the resultant force is increased while in the motion(in picture 2), hence the weight is increased (=apparent weight).

Notations

  • Please refer to the sketch attached

Calculations

  • Applying F=ma on the person down at the bottom towards the center of the circular path.

"\qquad
\begin{aligned}
\footnotesize R_1-mg&=\footnotesize \frac{mv^2}{r} \cdots\cdots\cdots 
R_1=\footnotesize mg+ \frac{mv^2}{r}\cdots(weight \,is\, increased\,=\,1.96\times10^3N)\\
\footnotesize v&=\footnotesize \sqrt{\frac{r(R_1-mg)}{m}}\\
& = \footnotesize \sqrt{ \frac{18m(1960N-655N)}{65.5kg}}\\
& =\footnotesize \bold {18.94 ms^{-1}}


\end{aligned}"

  • Velocity of the roller coaster equals that of the person riding it.